# 1]organization=Center for Research and Formation in Artificial Intelligence, Universidad de los Andes, Colombia, addressline=Carrera 1 No. 18a-12, city=Bogotá, postcode=111711, country=Colombia

### v1

@@ 32 VLMB [21] Mouse Brain 5 2 12202 100% 95% 29%
Intra- Patient 128 EHPCP2[5] Human Prostate cancer 10 1 24465 100% 92% 37%
@@ 3.2 Benchmark of Existing Gene Prediction Methods
We use SpaRED to evaluate eight state-of-the-art expression profile prediction methods. Among these, HisToGene Pang et al. (2021) splits a WSI into patches that are processed by a Visual Transformer (ViT) model. The output is the genetic profile of the WSI. STNet He et al. (2020) inputs individual patches into a fine-tuned DenseNet-121 with a linear layer for prediction. Additionally, STNet averages predictions across 8 symmetries of each patch to determine the final output. EGN Yang et al. (2023) applies exemplar-guided learning, a prediction strategy that bases its estimations on patches that are visually similar to the target patch within a latent space. This model integrates a ViT backbone with an Exemplar Bridging (EB) block, which dynamically improves feature representations using the most relevant exemplars. Building upon this approach, EGGN Yang et al. (2024) introduces an enhanced framework that, given a tissue slide image, encodes its windows into a feature space, retrieves exemplars from a reference dataset, constructs a graph, and dynamically predicts the gene expression of each window using an exemplar-guided graph network. BLEEP Xie et al. (2023) employs bi-modal contrastive learning to map image patches and expression profiles in a shared latent space, leveraging paired data to enhance representation learning. SEPAL Mejia et al. (2023) fine-tunes a ViT backbone and subsequently refines its predictions applying a GNN that processes a neighborhood graph for each patch. Additionally, SEPAL supervises expression changes relative to the mean expression in the training data instead of the absolute expression value, a strategy denoted ($\Delta$) prediction.
Furthermore, Hist2ST Zeng et al. (2022) first divides the input histology image into multiple patches, which are processed by a Convolutional Neural Network (CNN) to extract 2D visual features. These learned features are then passed through a Transformer, enabling the model to capture global dependencies within the WSI. The output is then processed by a Graph Neural Network (GNN) to capture spatial dependencies between neighboring patches. Finally, the resulting representations are used to predict gene expression levels. HGGEP Li et al. (2024) enhances feature extraction using a Gradient Enhancement Module (GEM). The latent features are then processed through a Convolutional Block Attention Module (CBAM) and a Vision Transformer (ViT), which leverage attention mechanisms to refine feature representations. A Hypergraph Association Module (HAM) further captures high-order associations by modeling global and local feature relationships based on spatial proximity.
@@
Remarkably, these two models rely on a graph-based approach that requires all spots in the WSI, resulting in high GPU memory consumption as the number of spots increases. Consequently, they can only run on 2 out of 26 SpaRED datasets. To evaluate these two models on the other 24 SpaRED datasets and thus enable a fair comparison with the other methods, we modify the model inputs by dividing the WSI into smaller sections (quarters, ninths, or sixteenths) and merging the predictions. To verify that this modification did not significantly impact the models’ performance, we conducted an experiment comparing the results of both models when using the same dataset, once with the complete WSI as input and once with the WSI divided into four parts. We used the VLO dataset for this experiment, as it is one of the datasets where the full image can be used as input. The results showed that dividing the image into four parts leads to an increase in MSE of 2.85% for Hist2ST and 0.002% for HGGEP, indicating a slight decrease in performance. These findings confirm that our methodology does not significantly affect the capacity of the model.
Alongside these models, our comprehensive benchmark also includes the performance of three baseline methods: a ShuffleNet Zhang et al. (2018) architecture that finetunes an image encoder with low computational cost, a ViT-B encoder Dosovitskiy et al. (2020) that reflects the impact of fine-tuning a state-of-the-art backbone for this task, and a ViT-B+$\Delta$ approach as suggested by Mejia et al. (2023). Moreover, we search for the optimal learning rate in every dataset. Then, with this value fixed, we explore two training scenarios: using raw data directly and SpaCKLE-completed data.
@@ 4 Data Completion with Transformers
Inspired by the disruptive success of the transformer architecture for completion tasks such as language next token prediction Vaswani et al. (2023) and visual reconstruction He et al. (2021), we adapt these ideas to the ST domain. Fig. 3 illustrates SpaCKLE’s training, a process that takes as a starting point data that we pre-completed using the median method proposed by Mejia et al. (2023). This process ensures faster training convergence and guarantees non-zero predictions.
Given the median-completed expression vector $x\in\mathbb{R}^{[g,1]}$ of a particular spot $s$ with $g$ prediction genes and the expression matrix $V_{x}\in\mathbb{R}^{[g,n]}$ that contains the genetic profile of the $n$ 2-hop neighbors in the Visium hexagonal geometry closest to $s$, we start by defining the expression matrix $E_{x}=[x;V_{x}]\in\mathbb{R}^{[g,n+1]}$. Knowing the neighborhood of spots around $s$, we also define matrix $M_{s}\in\{0,1\}^{[g,n+1]}$, a binary mask that presents through 0 values the gene expressions within each spot that were originally missing in the dataset and pre-completed using median values. Moreover, to implement the masked-autoencoder-like workflow, we construct the matrix $M_{rand}(\rho)\in\{0,1\}^{[g,n+1]}$, which randomly sets a fraction of $\rho=30\%$ values within the neighborhood as 1. The data points in $M_{rand}$ that have a value of 1 represent the elements in the neighborhood genetic profile that we set as candidates for our workflow to artificially hide. With these matrices, we then define the final random mask:
$M_{mask}=M_{s}\odot M_{rand}$ (1)
$M_{mask}$ determines which values to hide in our input data by setting a fraction of its elements to 0 as follows:
$E_{m}$ $=E_{x}\odot(1-M_{mask})$ (2)
Our purpose when combining these binary masks is to make sure that the data we randomly hide from the input expression matrix excludes values we pre-completed with the median expression of a given gene so that we do not consider the pre-completed values as ground truth when evaluating our model. Thus, we make sure that $M_{mask}$ does not overlap with median-completed points to ensure we have reliable ground truth values that were actually obtained with the ST technology when computing the evaluation metrics.
After randomly masking $E_{x}$, we process it with a transformer encoder $T(\cdot)$ that leverages the self-attention mechanism:
@@
$\mathcal{L}=\left\|E_{x}-\hat{E}_{x}\right\|_{2}^{2}$ (5)
However, we only compute metrics and complete missing values using the masked elements from the first vector of the output. Hence, each component of the completed version $\hat{x}$ can be expressed as:
$\hat{x}_{i}=\begin{cases}x_{i},&M_{mask}[1,i]=0\\ \hat{E}_{x}[1,i],&M_{mask}[1,i]=1\end{cases}$ (6)
@@ 4.1 Implementation Details:
We train all our models on a NVIDIA Quadro RTX 8000 with a batch size of 256 and use an Adam Kingma and Ba (2017) optimizer with default PyTorch library parameters. We handle both regression and completion problems as multivariate regression tasks and evaluate them using MSE and PCC. To select the best model, we save the one with the lowest validation MSE after 1,000 and 10,000 iterations for prediction and completion, respectively. All metrics are computed exclusively on real data for both the completion and the prediction task.
@@ 5.1 Gene Completion Evaluation
The violin plot in Fig. 4 presents a comparison of the logarithmic MSE for data completion using SpaCKLE, the median completion method, and stLearn across SpaRED. The results indicate that SpaCKLE outperforms alternative completion methods, with a relative 82.5% MSE reduction compared to the median method and by two orders of magnitude concerning stLearn. Notably, stLearn presents the highest MSE in the entirety of SpaRED, which conveys its inability to restore masked data. These results are consistent with those reported in Avşar and Pir (2023), where stLearn’s completion predictions included a high proportion of zero values. It is noteworthy that the median method is based solely on the adjacent expression of a single gene, an approach that, although straightforward, does not consider the broader genetic context. In contrast, SpaCKLE has access to the complete genetic profile of the neighboring spots. Thus, we hypothesize that our transformer architecture is leveraging the full expression profile of the empty spot’s vicinity to enhance completion predictions.
@@
Figure 5 presents additional examples of expression predictions made by SpaCKLE compared to those made by the median method for cases with 30% artificial missing data. The scatter plot visualizes the expression predictions against the ground truth across all the masked spots for a specific gene. Blue dots represent the predictions from SpaCKLE, while red dots correspond to the predictions from the median method. The black diagonal line indicates the ideal scenario in which the predictions perfectly match the ground truth. The results indicate that the blue dots follow the black line more closely compared to the red dots, suggesting that SpaCKLE has a greater capacity to accurately predict gene expression across most spots. The red dots, which represent the median method predictions, show a significant percentage of spots arranged in a straight line. This pattern indicates that many spots are assigned the same expression value. This observation supports the earlier analysis, which highlighted that the median method tends to predict a uniform value -corresponding to the global median- when it is unable to determine the local median.
@@
5.1.1 Ablation Experiments
To define the architecture of our final model, we explore variations that we hypothesized could potentially enhance SpaCKLE’s ability to complete missing data. For instance, in our first ablation experiment, we explore the use of visual information from the ST spots in the input neighborhood. In this case, the model’s workflow receives a matrix $H_{x}\in\mathbb{R}^{[d,n+1]}$ along with matrix $E_{m}$, described in detail in section 4. Matrix $H_{x}$ contains the image embeddings of dimension $d$ of the $n+1$ spots in the incoming neighborhood, which we obtain by processing their ST patches with a ViT model backbone that was fine-tuned for gene expression prediction, as proposed by Mejia et al. (2023). We concatenate the visual features with the genetic profile of its corresponding spot and feed this new matrix to the transformer encoder to continue with SpaCKLE’s framework.
In our second ablation experiment, we examine the effect of increasing the number of genes considered during completion. While SpaRED suggests using 128 or 32 genes per dataset, this parameter can be adjusted. To explore the impact of the length of the genetic profile, we define $x^{\prime}\in\mathbb{R}^{[1024,1]}$ as the extended median-completed expression vector containing 1024 genes, including the originally selected genes and additional ones with the next highest Moran’s I scores. In this way, we redefine the expression matrix as $E_{x^{\prime}}=[x^{\prime};V_{x^{\prime}}]\in\mathbb{R}^{[1024,n+1]}$. SpaCKLE processes and reconstructs the entire $E_{x^{\prime}}$ following the same masking and transformer encoding framework, incorporating an extended random mask $M^{\prime}_{rand}\in\{0,1\}^{[1024,n+1]}$. This mask ensures that values corresponding to the newly added genes remain zero, as they serve only as contextual information rather than targets for completion. By selectively masking only the predefined genes, we evaluate whether incorporating additional genetic context enhances the model’s predictive performance.
Table 2 shows the average performance of both alternative pipelines compared to the final SpaCKLE model’s performance when tested on all SpaRED datasets. These results show that including the embeddings of the histology patches does not enhance our model’s ability to complete missing values within a gene expression profile and actually increases its MSE by approximately 20% and decreases its PCC by 11%. Likewise, considering a wider set of genes for context represents an inconsistent enhancement of SpaCKLE’s performance, leading to an 10.36% increase in MSE and a 11.07% decrease in PCC. These results suggests that incorporating genes with weaker spatial autocorrelation provides no meaningful benefit for gene expression completion with SpaCKLE.
5.2 Gene Prediction Benchmark
Figure 6: Impact of SpaCKLE on SpaRED Benchmark. Prediction Pearson Correlation Coefficient for each model across all the datasets in SpaRED. For each dataset, the state-of-the-art model that obtains the highest Pearson Correlation Coefficient is included. As evidenced by the red (raw data) and blue dots (SpaCKLE-completed data), SpaCKLE improves performance across al methods in every dataset.
Table 2: Ablation Experiments Results. Comparison of SpaCKLE with alternative architectures incorporating visual features or additional context genes.
@@ 5.2 Gene Prediction Benchmark
SpaCKLE with Context Genes 0.787 0.534

### v2

@@ 32 VLMB [21] Mouse Brain 5 2 12202 100% 95% 29%
Intra-
Subject
128 EHPCP2[5] Human Prostate cancer 10 1 24465 100% 92% 37%
@@ 3.2 Benchmark of Existing Gene Prediction Methods
We use SpaRED to evaluate eight state-of-the-art expression profile prediction mds. Among these, STNet (He et al., 2020) inputs individual patches into a fine-tuned DenseNet-121 with a linear layer for prediction. Additionally, STNet averages predictions across 8 symmetries of each patch to determine the final output. HisToGene (Pang et al., 2021) splits a WSI into patches that are processed by a Visual Transformer (ViT) model. The output is the genetic profile of the WSI. Hist2ST (Zeng et al., 2022) divides the input histology image into multiple patches, which are processed by a Convolutional Neural Network (CNN) to extract 2D visual features. These learned features are then passed through a Transformer, enabling the model to capture global dependencies within the WSI. The output is then processed by a Graph Neural Network (GNN) to capture spatial dependencies between neighboring patches. Finally, the resulting representations are used to predict gene expression levels. BLEEP citepxie2023spatially employs bi-modal contrastive learning to map image patches and expression profiles in a shared latent space, leveraging paired data to enhance representation learning. SEPAL (Mejia et al., 2023) fine-tunes a ViT backbone and subsequently refines its predictions applying a GNN that processes a neighborhood graph for each patch. Additionally, SEPAL supervises expression changes relative to the mean expression in the training data instead of the absolute expression value, a strategy denoted ($\Delta$) prediction. EGN (Yang et al., 2023) applies exemplar-guided learning, a prediction strategy that bases its estimations on patches that are visually similar to the target patch within a latent space. This model integrates a ViT backbone with an Exemplar Bridging (EB) block, which dynamically improves feature representations using the most relevant exemplars. Building upon this approach, EGGN (Yang et al., 2024) introduces an enhanced framework that, given a tissue slide image, encodes its windows into a feature space, retrieves exemplars from a reference dataset, constructs a graph, and dynamically predicts the gene expression of each window using an exemplar-guided graph network. Finally, HGGEP (Li et al., 2024) enhances feature extraction using a Gradient Enhancement Module (GEM). The latent features are then processed through a Convolutional Block Attention Module (CBAM) and a Vision Transformer (ViT), which leverage attention mechanisms to refine feature representations. A Hypergraph Association Module (HAM) further captures high-order associations by modeling global and local feature relationships based on spatial proximity.
@@
Remarkably, Hist2ST and HGGEP rely on a graph-based approach that requires all spots in the WSI, resulting in high GPU memory consumption as the number of spots increases. Consequently, they can only run on 2 out of 26 SpaRED datasets. To evaluate these two models on the other 24 SpaRED datasets and thus enable a fair comparison with the other methods, we modify the model inputs by dividing the WSI into smaller sections (quarters, ninths, or sixteenths) and merging the predictions. To verify that this modification did not significantly impact the models’ performance, we conducted an experiment comparing the results of both models when using the same dataset, once with the complete WSI as input and once with the WSI divided into four parts. We used the VLO dataset for this experiment, as it is one of the datasets where the full image can be used as input. The results showed that dividing the image into four parts leads to an increase in MSE of 2.85% for Hist2ST and 0.002% for HGGEP, indicating a slight decrease in performance. These findings confirm that our methodology does not significantly affect the capacity of the model.
[…]
@@ 4 Data Completion with Transformers
Inspired by the disruptive success of the transformer architecture for completion tasks such as language next token prediction (Vaswani et al., 2023) and visual reconstruction (He et al., 2021), we adapt these ideas to the ST domain. Fig. 3 illustrates SpaCKLE’s training, a process that takes as a starting point data that we pre-completed using the median method proposed by Mejia et al. (2023). This process ensures faster training convergence, guarantees non-zero predictions, and improves the overall performance of our completion model, as demonstrated in Section 5.1.1.
Given the median-completed expression vector $x\in\mathbb{R}^{g}$ of a particular spot $s$ with $g$ prediction genes and the expression matrix $V_{x}\in\mathbb{R}^{g~\times~n}$ that contains the genetic profile of the $n$ 2-hop neighbors in the Visium hexagonal geometry closest to $s$, we start by defining the expression matrix $E_{x}=\begin{bmatrix}x&V_{x}\end{bmatrix}\in\mathbb{R}^{g~\times~(n+1)}$. Knowing the neighborhood of spots around $s$, we also define matrix $M_{s}\in\{0,1\}^{g~\times~(n+1)}$, a binary mask that presents through 0 values the gene expressions within each spot that were originally missing in the dataset and pre-completed using median values. Moreover, to implement the masked-autoencoder-like workflow, we construct the matrix $M_{rand}(\rho)\in\{0,1\}^{g~\times~(n+1)}$, which randomly sets a fraction of $\rho=30\%$ values within the neighborhood as 1. The data points in $M_{rand}$ that have a value of 1 represent the elements in the neighborhood genetic profile that we set as candidates for our workflow to artificially hide. With these matrices, we then define the final random mask as
$M_{mask}=M_{s}\odot M_{rand}.$ (1)
$M_{mask}$ determines which values to hide in our input data by setting a fraction of its elements to 0, as follows:
$E_{m}$ $=E_{x}\odot(1-M_{mask}).$ (2)
$M_{mask}$ is designed not to overlap with median-completed spots, guaranteeing that the ground truth used for computing evaluation metrics comes solely from values obtained with ST technology. After randomly masking $E_{x}$, we process it with a transformer encoder $T(\cdot)$ that leverages the self-attention mechanism:
@@
$\mathcal{L}=\left\|E_{x}-\hat{E}_{x}\right\|_{2}^{2}.$ (5)
However, we only compute metrics and complete missing values using the masked elements from the first vector of the output. Hence, each component of the completed version $\hat{x}$ can be expressed as
$\hat{x}_{i}=\begin{cases}x_{i},&M_{mask}[i,1]=0\\ \hat{E}_{x}[i,1],&M_{mask}[i,1]=1.\end{cases}$ (6)
@@ 4.1 Implementation Details:
We train all our models on a NVIDIA Quadro RTX 8000 with a batch size of 256 and use an Adam (Kingma and Ba, 2017) optimizer with default PyTorch library parameters. We train one completion model for each dataset, and optimize each one using a range of ten different learning rates sampled on a logarithmic scale between $1\times 10^{-5}$ and $1\times 10^{-2}$. We also conduct this learning rate optimization on the gene expression prediction models of our benchmark. Furthermore, we use a constant learning rate during training, and to ensure the reproducibility of our experiments, we fix the random seed to 42 and set all relevant random number generators accordingly.
We handle both regression and completion problems as multivariate regression tasks and evaluate them using MSE and PCC. To select the best model, we save the one with the lowest validation MSE after 1,000 and 10,000 iterations for prediction and completion, respectively. All metrics are computed exclusively on real data for both the completion and the prediction task.
@@ 5.1 Gene Completion Evaluation
The violin plot in Fig. 4 presents a comparison of the logarithmic MSE for data completion using SpaCKLE, the median completion method, and stLearn across SpaRED. The results indicate that SpaCKLE outperforms alternative completion methods, with a relative 82.5% MSE reduction compared to the median method and by two orders of magnitude concerning stLearn. Notably, stLearn presents the highest MSE in the entirety of SpaRED, which conveys its inability to restore masked data. These results are consistent with those reported in (Avşar and Pir, 2023), where stLearn’s completion predictions included a high proportion of zero values. It is noteworthy that the median method is based solely on the adjacent expression of a single gene, an approach that, although straightforward, does not consider the broader genetic context. In contrast, SpaCKLE has access to the complete genetic profile of the neighboring spots. Thus, we hypothesize that our transformer architecture is leveraging the full expression profile of the empty spot’s vicinity to enhance completion predictions.
@@
Fig. 5 presents additional examples of expression predictions made by SpaCKLE compared to those made by the median method for cases with 30% artificial missing data. The scatter plot visualizes the expression predictions against the ground truth across all the masked spots for a specific gene. Blue dots represent the predictions from SpaCKLE, while red dots correspond to the predictions from the median method. The black diagonal line indicates the ideal scenario in which the predictions perfectly match the ground truth. The results indicate that the blue dots follow the black line more closely compared to the red dots, suggesting that SpaCKLE has a greater capacity to accurately predict gene expression across most spots. The red dots, which represent the median method predictions, show a significant percentage of spots arranged in a straight line. This pattern indicates that many spots are assigned the same expression value. This observation supports the earlier analysis, which highlighted that the median method tends to predict a uniform value -corresponding to the global median- when it is unable to determine the local median.
@@
Table 2: Comparison of SpaCKLE and three ablated configurations: (i) using Context Genes, (ii) incorporating Visual Features, and (iii) training without data pre-completion. Metrics reported are average MSE and PCC on all SpaRED datasets. Best configuration is bolded.
@@
SpaCKLE without data pre-completion 6.166 0.165
Table 3: Effect of spatial neighborhood size on SpaCKLE’s completion metrics. Results for 0, 6, 18, and 36 neighbors (0–3 Visium hops) are shown, including average MSE and PCC on all SpaRED datasets. Best configuration is bolded.
Number of Neighbors
0 6 18 36
MSE 0.8136 0.7263 0.7134 0.7156
PCC 0.4727 0.5290 0.5316 0.5321
5.1.1 Ablation Experiments
To understand the contributions of each component in SpaCKLE, we carry out a series of controlled ablations on all SpaRED datasets. First, we assess the impact of our pre-completion step. In the full pipeline, we complete missing gene expression values with the median-completion method of Mejia et al. (2023) before training. In the ablated variant, we train directly on the raw data with missing entries. As shown in Table 2, median pre-completion reduces the mean squared error (MSE) by a factor of eight and increases the Pearson correlation coefficient (PCC) by nearly threefold, confirming that filling missing spots with a simple median estimate provides richer signals for learning and leads to substantially higher-quality reconstructions. To prevent the model from memorizing medians, we mask out only original, non-precompleted values during synthetic masking in training.
Next, we explore the use of visual information from the ST spots in the input neighborhood. In this case, the model’s workflow receives a matrix $H_{x}\in\mathbb{R}^{d~\times~(n+1)}$ along with matrix $E_{m}$, described in detail in section 4. Matrix $H_{x}$ contains the image embeddings of dimension $d$ of the $n+1$ spots in the incoming neighborhood, which we obtain by processing their ST patches with a ViT model backbone that we fine-tuned for gene expression prediction, as proposed by Mejia et al. (2023). We concatenate the visual features with the genetic profile of its corresponding spot and feed this combined representation into the transformer encoder as part of SpaCKLE’s framework.
In contrast to our original assumptions, the inclusion of visual features leads to a $\sim$20% increase in MSE and an $\sim$11% drop in PCC, indicating a reduction in completion performance (Table 2). These results prompt further analysis, and we suspect that several factors may explain this behavior. First, although we fine-tuned the ViT model for gene expression prediction, we freeze its weights during SpaCKLE’s training. This may have limited the capacity of the visual features to adapt to the masked reconstruction objective of our transformer-based framework. Additionally, we hypothesize that the use of domain-specific histology foundation models, such as UNI (Chen et al. (2024b)), could yield more task-relevant visual representations. Finally, another possible explanation is that the direct concatenation of gene expression vectors and visual embeddings, which are two modalities with inherently different distributions and scales, introduces imbalances that negatively impact learning. Our findings suggest that more sophisticated fusion mechanisms, such as modality-aware normalization, gated integration, or joint end-to-end training, may be beneficial for fully leveraging histological information in gene expression completion.
We also investigate the effect of profile length. While SpaRED defaults to selecting 32 or 128 genes by Moran’s I score, we extend this to 1024 genes, defining $x^{\prime}\in\mathbb{R}^{1024}$, $E_{x^{\prime}}=\begin{bmatrix}x^{{}^{\prime}}&V_{x^{\prime}}\end{bmatrix}\in\mathbb{R}^{1024~\times~(n+1)}$, where the extra genes have the next highest spatial autocorrelation. We mask only the original 32/128 genes using a random mask $M^{\prime}_{rand}\in\{0,1\}^{1024~\times~(n+1)}$ that ensures new genes remain zero for treating them purely as context. This wider context increases MSE by 10.4% and drops PCC by 11.1%, suggesting that genes with weaker spatial patterns add noise rather than helpful cues.
Finally, we examine how the number of spatial neighbors influences completion. We vary the neighborhood size among 0, 6, 18 and 36 spots - corresponding to 0, 1, 2, or 3 Visium hops, respectively - and retrained SpaCKLE with the same masking scheme. As Table 3 shows, the jump from 0 to 6 neighbors yields the largest improvement, expanding to two hops (18 neighbors) delivers a modest further improvement, but pushing to three hops (36 neighbors) produces a 0.3% increase in MSE and only 0.1% improvement in PCC. Given this plateau and the computational cost of larger neighborhoods, we select 18 neighbors as the optimal number of neighbors.
Together, these ablations confirm our design choices: (1) median pre-completion is essential for strong recovery, (2) a moderate neighborhood of 18 spots provides the best trade-off between performance and efficiency, and (3) neither adding histology embeddings nor context genes yields consistent benefit for SpaCKLE’s core completion task.
5.2 Gene Prediction Benchmark
Figure 6: Impact of SpaCKLE on SpaRED Benchmark. Prediction Pearson Correlation Coefficient for each model across all the datasets in SpaRED. For each dataset, the state-of-the-art model that obtains the highest Pearson Correlation Coefficient is included. As evidenced by the red (raw data) and blue dots (SpaCKLE-completed data), SpaCKLE improves performance across all methods in every dataset.